The demand function for watches is: Q(p)=160-4p. Calculate the marginal revenue at output level (q) equal to 20.
Answer #1
Q = 160-4P
For Q = 20
20 = 160 - 4P
P = 140 / 4 = 35
Total revenue 160P - 4P^2
Marginal revenue 160 - 8P
Marginal revenue for Q=20, P=35
160 - 8x35 = -120